Welland Co. shows the following information on its 2016 income statement: sales $155,500; costs = S81,400, other expenses-$4,700, depreciation expense $10,400; interest expense $7,900; taxes = $17,885; dividends = $7,450. In addition, youre told that the frm issued $3.100 in new equity during 2016 and redeemed $4,900 in outstanding long-term debt. a. What is the 2016 operating cash flow? (Do not round Intermedlate calculations and round your answer to the nearest whole number, e.g 32.) Operating cash flow 51513 b. What is the 2016 cash flow to creditors? (Do not round Intermediate calculations and round your answer to the nearest whole number, e.g 32.) Cash flow to creditorsS c. what is the 2016 cash flow to stockholders? (Do not round intermediate calculations and round your answer to the nearest whole number, .g,32.) Cash flow to stockholders $ d. If net fixed assets Iincreased by $20,500 during the year, what was the addition to NWC? (Do not round intermediate calculations and round your answer to the nearest whole number, e-g,32.) Addition to net working capital
Answer #1

a) Operating Cash Flow = Sales - Costs - Other Expenses -Taxes = 155,500 - 81,400 - 4,700 - 17,885 = 51515

b) Cash Flow to creditors = Dividends - New Equity = 7,450 - 3100 = 4350

c) Cash flow to debt holders = Interest - New Debt = 7900 + 4900 = 12,800

d) Cash Flow from Assets = Cash Flow to creditors + Cash Flow to Debt Holders = 4350 + 12,800 = 17150
Cash Flow from Assets = OCF - increase in fixed assets - increase in Working capital
17150 = 51515 - 20,500 - Increase in Working capital
Increase in Working capital = 51515 - 20500 - 17150 = 3465

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well